# Not All LLM-Generated Data Are Equal: Rethinking data weighting in Text Classification

### v1

@@ 5 Experiments
Table 1: Performance metrics across datasets and methods. The table showcases the accuracy (Acc) and macro F1 score (F1) for each method applied on three benchmarks. The methods include GPT-3.5 few-shot, CE-Loss, Focal-Loss, Hu et al.’s method, IMP-Loss, and DIMP-Loss. Bold entries denote the performance within 0.5%, comparing to the best performance of each training source.
We assessed our proposed methods by comparing them with standard loss functions across several text classification benchmarks, including Financial Phrasebank (Financial) (Malo et al., 2014), irony detection (Tweet Irony) (Van Hee et al., 2018), and the MRPC dataset from GLUE (Wang et al., 2018). Detailed descriptions and specifications of these datasets are provided in Appendix H. In our experiments, we referred the large real-world data $D_{P}$ to the original training set from each benchmark and the small real-world data $D_{P^{\prime}}$ to the original development set, with the sizes from approximately 200 to 400, as shown in Table 5. Our experiments explored three different scenarios: training solely on synthetic data (Sec. 5.1), real-world data (Sec. 5.2), and noisy data (Sec. G). We evaluated Accuracy (Acc) and Macro F1 score (F1) for every benchmark. These metrics were computed by comparing the model’s predictions with the gold labels provided in the test sets. We used a BERT-based model for fine-tuning and building the checkers. The Appendix F details the configurations.
@@
CE-Loss, Focal Loss, and Hu et al. (2019)’s method are our baselines, detailed in Sec. 2.1 and Sec. 2.2. Focal Loss addressed class imbalance and mitigated easily classified data’s impact, preventing overconfidence. The weight function for Focal Loss was defined as $w_{i}=(1-\hat{P}(y_{i}|\mathbf{x}_{i};\theta))^{\gamma}$ where $\gamma$ controled the downweighting of easy examples. Mukhoti et al. (2020) showed models trained with Focal Loss exhibited better calibration under the i.i.d. assumption and performed robustly under distribution shifts. This made Focal Loss a strong baseline for evaluating our proposed approaches. Hu et al. (2019) used meta-learning to dynamically adjust weights training data by maximizing the likelihood on a small real-world dataset, similar to IMP-Loss and DIMP-Loss; however, our methods directly adjust weights based on quality and diversity checkers, while Hu et al. (2019)’s method relied on meta-learning to optimize weights indirectly.  [^2] We implemented Focal-Loss and Hu et al. (2019)’s method by using their official code.
@@ 5.1 Training on LLM-generated Data
We used GPT-3.5-turbo-1106 (OpenAI, 2022). A system prompt, 8-shot examples from the development set $D_{P^{\prime}}$, and the corresponding labels to generate the input text were provided. For the Financial and Tweet Irony benchmarks, our generation prompt based on previous research (Li et al., 2023). Similarly, for the MRPC benchmark, the prompt included pairs of sentences with answers, which automatically guided the LLM in generating the answers. The Appendix I.1 provides each benchmark’s prompt.
@@
As shown in Table 1, our methods outperformed all baselines in all benchmarks. For instance, in Financial Phrasebank, IMP-Loss achieved 82.09% in accuracy and 79.40% in F1 score, and DIMP-Loss reached 82.67% and 79.53% respectively, while the CE-Loss reached 77.39% / 74.01%. The result showed if we use the baselines, CE-Loss, Focal-Loss, and Hu et al. (2019)’s method to train a classifier on $D_{Q}$, the performance could be worse than that of using CE-Loss on much smaller $D_{P^{\prime}}$ (quality checker). In contrast, our methods consistently outperformed the small quality checker, encouraging the usage of abundant LLM-generated data. Notably, even when the quality checker performs poorly, such as on the Tweet Irony dataset, where the accuracy was 62.5%, which was lower than the 76.9% achieved by directly training on generated data using CE-Loss, our methods still delivered strong performance. This suggested that a high-performance quality checker was not a prerequisite for the effectiveness of our methods.
@@
Figure 1: Training dynamics showing the testing accuracy over five epochs for the three benchmarks. This chart displays the minimum, maximum, and average accuracy observed across four runs with different random seeds, comparing our proposed methods with the standard CE-Loss and Focal-Loss.
@@ 5.2 Training on Real World Data
As shown in Table 1, IMP-Loss and DIMP-Loss outperformed other baselines even when applied directly to real-world data. Although the performance improvements are less than that of using GPT-3.5-generated data, the results indicated our methods were versatile and able to handle multiple sources of training data effectively. Specifically, in the Financial dataset, IMP-Loss achieved 85.3% Acc and 83.27% F1 score, while DIMP-Loss reached 85.4% Acc and 82.79% F1 score, surpassing both CE-Loss, Focal-Loss, and Hu et al.’s method baselines. From our perspective, the reduced improvement in this scenario was due to the lack of a requirement to shift the training data distribution. Regarding the asymptotic viewpoint, the optimal solution of cross-entropy is already the best solution when training on real-world data. Nonetheless, our methods demonstrated robust performance across various conditions.
Figure 2: Test accuracy on the Financial benchmark with varying percentages of the training set for the quality checker. The graph shows the performance of IMP-Loss, DIMP-Loss, and their Quality Checker.
@@ 6 Related Works
We list some essential related work in this section and others in A.
Synthetic Data Generation from LLMs.
Recent advancements in generative AI have spurred interest in using LLMs to generate synthetic data for specific tasks, particularly in low-resource settings. Studies have explored zero-shot and few-shot settings for data generation, where LLMs directly generate instances with distinct labels or use little real-world data as examples to create relevant and diverse data (Li et al., 2023; West et al., 2022). LLMs have the unique ability to generate both labels and diverse input instances, significantly enhancing the variety and quality of synthetic datasets. Approaches like ZEROGEN synthesized data by pre-trained LMs to train smaller task models, achieving competitive performance in NLP tasks (Ye et al., 2022). Additionally, (Wang et al., 2023; Taori et al., 2023) have fine-tuned GPT-3 and LLaMA on augmented data, respectively, to follow human instructions, while domain-specific tuning before generation has been effective for tasks like text classification (Stylianou et al., 2023).
@@
Importance weighting (IW) serves as a classical strategy for addressing the issue of shifts between data distributions (Hesterberg, 1995). In traditional applications of IW, weights are derived by evaluating the degree of similarity between training and testing distributions through various statistical techniques. These techniques include maximum mean discrepancy (Schölkopf et al., 2007) and the estimation of KL divergence (Sugiyama et al., 2007). Although effective in linear model contexts, the efficacy of these methods seems to significantly diminish when applying IW to more complex deep learning frameworks (Byrd & Lipton, 2019). Besides traditional methods, recent studies have explored approaches such as Focal Loss (Lin et al., 2017) and meta-learning techniques (Hu et al., 2019), which address issues like class imbalance and dynamic reweighting, as discussed in Sec. 2.2.
@@ 7 Conclusions and Discussions
IMP-Loss and DIMP-Loss are novel weighted-loss objectives that further enhance the performance of models trained on LLM-generated data. Our empirical results demonstrated that both methods outperformed traditional loss functions across various benchmarks. Notably, DIMP-Loss was particularly computationally efficient, requiring subtly additional resources while increasing performance. These findings emphasized the potential of IMP-Loss and DIMP-Loss in effectively leveraging synthetic data for training machine learning models. In the future, we will extend our methods on question answering, text generation, LLM pertaining, and other potential tasks, further exploring how quality and diversity matter for learning.
@@ 8 Reproducibility Statement
To ensure reproducibility, we provided the source code and generated dataset in supplementary materials, prompts used for generation in I.1, hyper-parameters and other training details in F, testing datasets’ descriptions in H, and theoretical results in B, C, D. In addition, for the baselines, we implemented the Focal Loss as in the source code and used the publicly available code on (Hu et al., 2019)’s Github but replaced the input data. We hope one can smoothly reproduce our results via these materials.
@@ Appendix C Derivation of DIMP-Loss
In this section, we provide the formal derivation to address the question: Which data point in $D_{Q}$, when used for training, will most effectively bring the model distribution closest to $P$? Following the optimization formulation in Eq. 7, we can empirically apply Monte Carlo estimation using a small real-world dataset $D_{P^{\prime}}$, denoted as $(\mathbf{y}_{P^{\prime}},\mathbf{x}_{P^{\prime}})$. This allows us to reformulate the problem by maximizing the joint probability of the data points in $D_{P^{\prime}}$, which leads to the following optimization problem. This derivation is similar to the online batch selection techniques discussed in previous research (Deng et al., 2023)
$\operatorname*{arg\,max}_{(\mathbf{x},y)\in D_{Q}}\hat{P}(\mathbf{y}_{P^{\prime}}|\mathbf{x}_{P^{\prime}};\theta_{t},\{(\mathbf{x},y)\})$ (21)
@@
$=\frac{\hat{P}(D_{P^{\prime}},\mathbf{x},y,\theta_{t})}{\hat{P}(\mathbf{x}_{P},\mathbf{x},y,\theta_{t})}$
@@ Appendix E Computational Time Comparison
Table 3: Total running time of each component (in seconds) for CE-Loss, IMP-Loss, and DIMP-Loss on the LLM-generated Financial benchmark. The table breaks down the time spent building the Quality Checker (QC), building the Diversity Checker (DC), precalculating weights, and training. The total time combines all these components.
In this computational time experiment, we evaluated the running times on the LLM-generated Financial benchmark dataset, which includes 10,012 training samples ($D_{Q}$) and 242 development samples (small real-world data, $D_{P^{\prime}}$). Our comparison focused on three methods: CE-Loss, IMP-Loss, and DIMP-Loss. We have broken down the total process into four components: building the Quality Checker (QC), building the Diversity Checker (DC), precalculating constant weights, and the actual training time for each respective loss function. For all experiments, the downstream models and checkers were trained for 5 epochs with a batch size of 32. The batch size was set to 64 during the precalculating constant weights phase. The results of this breakdown are presented in Table 3, and the total time is in Figure 3. The results indicate that IMP-Loss requires approximately 2.2 times the running time of CE-Loss. In contrast, DIMP-Loss requires only about 1.1 times the running time of CE-Loss. This demonstrates that training with DIMP-Loss is highly efficient, adding only a slight overhead while significantly enhancing downstream performance.
@@ Appendix G Training on noisy data
Noisy Data CE-Loss 79.95 77.41 60.46 60.14 74.03 67.5
Focal-Loss 80.2 78.26 62.11 61.12 74.72 69.59
IMP-Loss (Ours) 81.85 78.93 64.8 64.51 76 70.46
DIMP-Loss (Ours) 82.51 80.22 64.16 64.09 76.58 71.32
@@
In this section, we evaluate the robustness of our proposed methods, IMP-Loss and DIMP-Loss, by training on noisy datasets. We aim to simulate real-world scenarios where LLM-generated data may be imperfect due to labeling errors (low quality) or duplicate entries (low diversity). This allows us to analyze the effects of the Quality Checker and Diversity Checker in IMP-Loss and DIMP-Loss.
@@ G.2 Performance Results
The results in Tabel 4 indicated that our proposed methods, IMP-Loss and DIMP-Loss, consistently outperform the baselines across all benchmarks, even when the training data is noisy. Specifically, in the Financial dataset, IMP-Loss achieves 81.85% Acc and 78.93% F1 score, while DIMP-Loss reaches 82.51% Acc and 80.22% F1 score, surpassing the CE-Loss and Focal-Loss baselines. In the Tweet Irony dataset, the performance improvement is more pronounced, with IMP-Loss achieving 64.8% Acc and 64.51% F1 score, and DIMP-Loss achieving 64.16% Acc and 64.09% F1 score, significantly higher than CE-Loss and Focal-Loss. For the MRPC dataset, IMP-Loss and DIMP-Loss show robust performance with 76% Acc and 70.46% F1 score and 76.58% Acc and 71.32% F1 score, respectively, outperforming the GPT-3.5 few-shot approach, which achieves 69.28% Acc and 71.75% F1 score.

### v2

@@ 5 Experiments
Table 1: Performance metrics across datasets and methods. The table showcases each combination’s accuracy (Acc) and macro F1 score (F1). The methods include GPT-3.5 few-shot, CE-Loss, Focal-Loss, Hu et al.’s method, SunGen, IMP-Loss, and DIMP-Loss. Bold entries denote the performance within 0.5%, compared to the best performance of each training source.
We assessed our proposed methods by comparing them with standard loss functions across several text classification benchmarks, including Financial Phrasebank (Financial) (Malo et al., 2014), irony detection (Tweet Irony) (Van Hee et al., 2018), and the MRPC dataset from GLUE (Wang et al., 2018). Detailed descriptions and specifications are provided in Appendix H. In our experiments, we referred the large real-world data $D_{P}$ to the original training set from each benchmark and the small real-world data $D_{P^{\prime}}$ to the original development set, with the sizes from approximately 200 to 400, as shown in Table 5. Our experiments explored three different scenarios: training solely on synthetic data (Sec. 5.1), real-world data (Sec. 5.2), and noisy data (Sec. G). We evaluated Accuracy (Acc) and Macro F1 score (F1) for every benchmark. These metrics were computed by comparing the model’s predictions with the gold labels provided in the test sets. We used a BERT-based model for fine-tuning and building the checkers. The Appendix F details the configurations.
@@
CE-Loss, Focal Loss, SunGen and Hu et al. (2019) are our baselines, detailed in Sec. 2.1 and Sec. 2.2. Focal Loss addressed class imbalance and mitigated easily classified data’s impact, preventing overconfidence. The weight function for Focal Loss was defined as $w_{i}=(1-\hat{P}(y_{i}|\mathbf{x}_{i};\theta))^{\gamma}$ where $\gamma$ controled the downweighting of easy examples. Mukhoti et al. (2020) showed models trained with Focal Loss exhibited better calibration under the i.i.d. assumption and performed robustly under distribution shifts. This made Focal Loss a strong baseline for evaluating our proposed approaches. Both SunGen and Hu et al. (2019) are bilevel optimization methods but differ in their weight update mechanisms and the objective functions of their outer loops. They used meta-learning to dynamically adjust weights training data by maximizing the likelihood on a small real-world dataset, similar to IMP-Loss and DIMP-Loss; however, our methods directly adjust weights based on quality and diversity checkers, while Hu et al.’s method and SunGen relied on meta-learning to optimize weights indirectly  [^2] We implemented Focal-Loss, SunGen and Hu et al. (2019) by using their official code..
@@ 5.1 Training on LLM-generated Data
We used GPT-3.5-turbo-1106 (OpenAI, 2022), given a system prompt, 8-shot examples from the development set $D_{P^{\prime}}$, and the corresponding labels to generate the input text. For the Financial and Tweet Irony, our generation prompt based on previous research (Li et al., 2023). Similarly, for the MRPC benchmark, the prompt included pairs of sentences with answers, which automatically guided the LLM in generating the answers. See Appendix I.1 for details.
@@
As shown in Table 1, our methods outperformed all baselines in all benchmarks. For instance, in Financial Phrasebank, IMP-Loss achieved 82.09% in accuracy and 79.40% in F1 score, and DIMP-Loss reached 82.67% and 79.53% respectively, while the CE-Loss reached 77.39% / 74.01%. The result showed if we use the baselines, CE-Loss, Focal-Loss, SunGen and Hu et al. (2019) to train a classifier on $D_{Q}$, the performance could be worse than that of using CE-Loss on much smaller $D_{P^{\prime}}$ (quality checker). In contrast, our methods consistently outperformed the small quality checker, encouraging the usage of abundant LLM-generated data. Notably, even when the quality checker performs poorly, such as on the Tweet Irony dataset, where the accuracy was 62.5%, which was lower than the 76.9% achieved by directly training on generated data using CE-Loss, our methods still delivered strong performance. This suggested that a high-performance quality checker was not a prerequisite for the effectiveness of our methods. Although the performance of the meta-learning-based SunGen method was, in some cases, close to that of our methods (though still slightly below), our approaches have significant advantages in computational efficiency, making them more practical for large-scale applications. Further details on computational efficiency are shown in the Appendix E.
@@
Figure 1: Training dynamics shows the testing accuracy over five epochs for benchmarks. This chart displays the minimum, maximum, and average accuracy observed across four runs with different random seeds, comparing our proposed methods with the standard CE-Loss and Focal-Loss.
@@ 5.2 Training on Real World Data
As shown in Table 1, IMP-Loss and DIMP-Loss outperformed other baselines even when applied directly to real-world data. Although the performance improvements are less than that of using GPT-3.5-generated data, the results indicated our methods were versatile and able to handle multiple sources of training data effectively. Specifically, in the Financial dataset, IMP-Loss achieved 85.3% Acc and 83.27% F1 score, while DIMP-Loss reached 85.4% Acc and 82.79% F1 score, surpassing CE-Loss, Focal-Loss, and (Hu et al., 2019). From our perspective, the reduced improvement in this scenario was due to the lack of a requirement to shift the training data distribution. Regarding the asymptotic viewpoint, the optimal solution of cross-entropy is already the best solution when training on real-world data. Nonetheless, our methods demonstrated robust performance across various conditions.
Figure 2: Test accuracy on the Financial with varying percentages of the training set for the quality checker. The graph shows the performance of each loss and the Quality Checker.
@@ 6 Related Works
We list some essential related works in this section and others in A.
@@
Importance weighting (IW) serves as a classical strategy for addressing the issue of shifts between data distributions (Hesterberg, 1995). In traditional applications of IW, weights are derived by evaluating the degree of similarity between training and testing distributions through various statistical techniques. These techniques include maximum mean discrepancy (Schölkopf et al., 2007) and the estimation of KL divergence (Sugiyama et al., 2007). Although effective in linear model contexts, the efficacy of these methods seems to significantly diminish when applying IW to more complex deep learning frameworks (Byrd & Lipton, 2019). Besides traditional methods, recent studies have explored approaches such as Focal Loss (Lin et al., 2017) and meta-learning techniques (Hu et al., 2019; Meng et al., 2023; Gao et al., 2023), which take the weights of samples as trainable hyperparameters as discussed in Sec. 2.2.
Synthetic Data Generation from LMs
Recent advancements in generative AI have spurred interest in using LMs to generate synthetic data for specific tasks, particularly in low-resource settings. Studies have explored zero-shot and few-shot settings for data generation, where LMs directly generate instances with distinct labels or use little real-world data as examples to create relevant and diverse data (Li et al., 2023; West et al., 2022; Ye et al., 2022; Wang et al., 2023; Taori et al., 2023). LMs have the unique ability to generate both labels and diverse input instances, significantly enhancing the variety and quality of synthetic datasets. Approaches like ZEROGEN synthesized data by pre-trained LMs to train smaller task models, achieving competitive performance in NLP tasks, such as text classification (Ye et al., 2022).
LM-Generated Data for Training Text Classifier
Several studies have investigated leveraging LM-generated data for text classification. Some works maintain data quality by filtering strategy (Stylianou et al., 2023; Meng et al., 2022; 2023; Li et al., 2023; West et al., 2022; Ye et al., 2022). Another common approach is data reweighting. For example, SunGen (Gao et al., 2023) adopted a bilevel optimization approach to learn weights for synthetic data, incorporating a noise-robust loss in the outer loop to improve the reliability, and this benefited SunGen to outperform counterparts using meta-learning for data reweighting, such as Hu et al. (2019). Despite its advantages, the bilevel optimization process remains computationally expensive, making our approaches outstand by their efficiency. Moreover, there exists novel research further enhancing the use of synthetic data. For instance, UniGen (Choi et al., 2024) utilized contrastive learning to improve generalization capabilities, but required a open-source pretrained LM, while FuseGen (Zou et al., 2024) combined synthetic data from multiple LLMs to enhance performance. It is worth noting our approaches are compatible with UniGen or FuseGen, and a potential complement and enhancement of these works.
@@ 7 Conclusions and Discussions
IMP-Loss and DIMP-Loss are novel weighted-loss objectives that further enhance the performance of models trained on LLM-generated data. Our empirical results demonstrated that both methods outperformed traditional loss functions across various benchmarks. Notably, DIMP-Loss was particularly computationally efficient, requiring subtly additional resources while increasing performance. These findings emphasized the potential of IMP-Loss and DIMP-Loss in effectively leveraging synthetic data for training machine learning models. In the future, we will extend our methods on question answering, text generation, LLM pretraining, and other potential tasks, further exploring how quality and diversity matter for learning.
@@ 8 Reproducibility Statement
To ensure reproducibility, we provided the source code and generated dataset in supplementary materials, prompts used for generation in I.1, hyper-parameters and other training details in F, testing datasets’ descriptions in H, and theoretical results in B, C, D. In addition, for the baselines, we implemented the Focal Loss as in the source code and used the publicly available code provided by Hu et al. (2019) but replaced the input data. We hope one can smoothly reproduce our results via these materials.
Acknowledgments
We sincerely appreciate the insightful and valuable feedback from the anonymous reviewers. This work is primarily supported by the National Science and Technology Council, Taiwan, under grant number NSTC112-2221-E-001-025 and partially supported by E.SUN COMMERCIAL BANK. We extend our special thanks to Chien-An Chen and Yi-Ren Yeh at E.SUN for their valuable input during the preliminary discussions. Additionally, we thank the National Center for High-performance Computing (NCHC), Taiwan, for providing essential computational and storage resources.
@@ Appendix C Derivation of DIMP-Loss
In this section, we provide the formal derivation to address the question: Which data point in $D_{Q}$, when used for training, will most effectively bring the model distribution closest to $P$? Following the optimization formulation in Eq. 7, we can empirically apply Monte Carlo estimation using a small real-world dataset $D_{P^{\prime}}$, denoted as $(\mathbf{y}_{P^{\prime}},\mathbf{X}_{P^{\prime}})$. This allows us to reformulate the problem by maximizing the joint probability of the data points in $D_{P^{\prime}}$, which leads to the following optimization problem. This derivation is similar to the online batch selection techniques discussed in previous research (Deng et al., 2023).
$\operatorname*{arg\,max}_{(\mathbf{x},y)\in D_{Q}}\hat{P}(\mathbf{y}_{P^{\prime}}|\mathbf{X}_{P^{\prime}};\theta_{t},\{(\mathbf{x},y)\})=\\ \operatorname*{arg\,max}_{(\mathbf{x},y)\in D_{Q}}\prod_{(\mathbf{x}^{\prime},y^{\prime})\in D_{P^{\prime}}}\hat{P}(y^{\prime}|\mathbf{x}^{\prime};\theta_{t},\{(\mathbf{x},y)\})$ (21)
@@
$=\frac{\hat{P}(D_{P^{\prime}},\mathbf{x},y,\theta_{t})}{\hat{P}(\mathbf{X}_{P^{\prime}},\mathbf{x},y,\theta_{t})}$
@@ Appendix E Computational Time Comparison
Table 3: Total running time of each component (in seconds) for CE-Loss, SunGen, IMP-Loss, and DIMP-Loss on the LLM-generated financial benchmark. The table breaks down the time spent building the Quality Checker (QC), building the Diversity Checker (DC), precalculating weights, and training. The total time combines all these components.
In this computational time experiment, we evaluated the running times on the LLM-generated Financial benchmark dataset, which includes 10,012 training samples ($D_{Q}$) and 242 development samples (small real-world data, $D_{P^{\prime}}$). Our comparison focused on four methods: CE-Loss, SunGen, IMP-Loss, and DIMP-Loss. We have broken down the total process into four components: building the Quality Checker (QC), building the Diversity Checker (DC), precalculating constant weights, and the actual training time for each respective loss function. For all experiments, the downstream models and checkers were trained for 5 epochs with a batch size of 32. The batch size was set to 64 during the precalculating constant weights phase. The inner loop epochs were set to 1 for SunGen. The results of this breakdown are presented in Table 3, and the total time is in Figure 3. The results indicate that IMP-Loss requires approximately 2.2 times the running time of CE-Loss. In contrast, This demonstrates that DIMP-Loss is highly efficient, requiring only a slight overhead compared to CE-Loss, while SunGen’s computational time is approximately 7 times higher, further underscoring the efficiency of our methods for large-scale applications.
@@ Appendix G Training on noisy data
Noisy Data CE-Loss 78.38 73.44 60.46 60.14 74.03 67.5
Focal-Loss 78.55 74.97 62.11 61.12 74.72 69.59
IMP-Loss (Ours) 81.6 78.24 64.8 64.51 76 70.46
DIMP-Loss (Ours) 82.59 80.28 64.16 64.09 76.58 71.32
@@
In this section, we evaluate the robustness of our proposed methods, IMP-Loss and DIMP-Loss, by training on noisy datasets. We aim to simulate real-world scenarios where LLM-generated data may be imperfect due to labeling errors (low quality), duplicate entries (low diversity), and unrelated inputs (low quality). This allows us to analyze the effects of the Quality Checker and Diversity Checker in IMP-Loss and DIMP-Loss.
@@ G.1 Experimental Setup
  4. 4.
    Unrelated Input Data (Only for Financial): For the financial benchmark, we introduce out-of-domain input noise by randomly selecting 452 data points from the Tweet Sentiment Extraction benchmark (Maggie et al., 2020). While this dataset is also a sentiment classification task, it is unrelated to the financial domain.
@@ G.2 Performance Results
The results in Tabel 4 indicated that our proposed methods, IMP-Loss and DIMP-Loss, consistently outperform the baselines across all benchmarks, even when the training data is noisy. Specifically, in the Financial dataset, IMP-Loss achieves 81.6% Acc and 78.24% F1 score, while DIMP-Loss reaches 82.59% Acc and 80.28% F1 score, surpassing the CE-Loss and Focal-Loss baselines. In the Tweet Irony dataset, the performance improvement is more pronounced, with IMP-Loss achieving 64.8% Acc and 64.51% F1 score, and DIMP-Loss achieving 64.16% Acc and 64.09% F1 score, significantly higher than CE-Loss and Focal-Loss. For the MRPC dataset, IMP-Loss and DIMP-Loss show robust performance with 76% Acc and 70.46% F1 score and 76.58% Acc and 71.32% F1 score, respectively, outperforming the GPT-3.5 few-shot approach, which achieves 69.28% Acc and 71.75% F1 score.